# Predict. Optimize. Revise. On Forecast and Policy Stability in Energy Management Systems

### v3

@@ 1 Introduction
To effectively navigate these challenges, it is beneficial to integrate forecasting of the problem environment with the optimization of decision-making processes. This integration, often referred to as the predict, then optimize approach, is prevalent in energy applications like electric vehicle charging [1, 2], battery scheduling [3, 4], and energy and flexibility dispatch problems [5, 6]. In many of these applications, decision-makers face switching costs — expenses incurred from updating operational plans. For example, transitioning operations from one state to another often involves costs due to physical limitations on ramping energy production up or down [7], managing server load balancing [8], or covering network fees associated with energy trading [9].
Despite extensive research on online optimization with predictions, there remains a notable gap concerning how switching costs affect the synergy between forecasting models and optimization algorithms. This study addresses this gap by exploring the concepts of time-coupling and forecast stability within such systems. Time-coupling, where one decision directly affects subsequent decisions and system states, gives rise to switching costs. Forecast inaccuracies often lead to policy revisions, requiring adjustments to better align with the current system state and updated information. Striking a balance between forecast accuracy and the impact of switching costs is therefore crucial. This paper explores how this balance shapes system design and informs operational strategies.
Building on this interplay between forecast accuracy and switching costs, we propose that the stability of forecasts plays an equally critical role. Since forecasts in optimization are typically updated with each new data acquisition—covering multiple future steps and undergoing regular revisions—stability becomes essential. A forecast that remains consistent across updates reduces the need for frequent strategic adjustments, thereby mitigating the impact of switching costs. Consequently, this research aims to answer the following questions:
@@ 1.1 Related work
The third category, Semi-direct methods, considers characteristics of the optimization problem but abstains from direct interaction during training of the forecast. These methods can take various forms, as potential improvements can occur at any stage outside of the training process. An example is found in the work by Kazmi and Paskevich [13], where those authors propose a Bayesian Optimization approach for integrating the downstream optimization problem into the hyperparameter tuning process of the forecast model. Other example is [3], where the authors propose iterating attention over training data statically and dynamically to improve the forecast model performance in the downstream task.
There is a noticeable gap in the literature when it comes to studying operational aspects of forecasting model and optimization deployment. For forecasting, the frequency of the forecast revision can be interpreted as both the frequency of the forecast model deployment and the frequency of model training. The latter aspect is studied by Spiliotis and Petropoulos [14], where those authors investigate different scenarios of updating the model fit for univariate exponential smoothing (ES) and univariate gradient boosting models (LightGBM).
We are particularly interested in the applications of optimization with predictions in environments where switching costs are present. This topic has featured in research on online convex optimization (OCO) problems. In control theory, a widespread strategy for tackling online multi-step optimization challenges is the employment of the Receding Horizon Control (RHC) algorithm. Model Predictive Control (MPC) is the most common implementation of RHC that employs a dynamic model of the process and is widely used in online optimization. When rerun at every time step, this method commits to the first step of the future horizon while treating later decisions as advisory. The Fixed Horizon Control (FHC) algorithm is a generalization of the algorithm that commits to a fixed number of steps $v$ in the future before re-optimizing the policy. The difference between the two algorithms is illustrated in Figure 1. Throughout this paper, we will use the term FHC to align with the focus on commitment-specific optimization.
@@
Forecast stability has been recognized as a critical factor in planning. Building on the comprehensive framework by Godahewa et al. [18], which categorizes stability into vertical (variations across forecast updates for the same period) and horizontal (variations within a prediction window), this work investigates these dimensions in the context of energy management systems. Unlike supply chain management, where freezing intervals [19, 20] and smoothing techniques [21] have been widely used to address system nervousness and the bullwhip effect [22], energy systems require more dynamic and frequent forecast updates.
Nevertheless, switching costs are observed in energy management systems, such as ramping costs for generating units [7, 23] or toggling server load balancing [8, 24]. However, the literature lacks a systematic exploration of how commitment levels influence the interplay between forecast stability and switching costs. By focusing on these aspects, the current research sets the stage for advancing the design and operation of energy management systems, particularly in applications involving energy storage optimization and real-time decision-making.
@@ 2.1 Problem statement
This paper addresses the challenge of optimizing battery storage management within a multi-agent energy management system. The setting of the problem is first outlined in the CityLearn Challenge 2022 [25] and based on a real-world dataset from a grid-interactive community. The setup of the problem is shown in the Figure 2. More generally, we consider an online optimization with switching costs. The problem is solved with the fixed horizon control (FHC), the variation of model predictive control (MPC) with a fixed commitment horizon. The full details of the objective score and relevant constraints are provided in A.1. The problem includes finding an optimal state-dependent policy $X_{t}^{FHC}(S_{t}|\theta^{LA})$. The decision-maker is faced with a sequence of decisions, where the decision at each step is based on the current state $S_{t}$ of the system and the direct lookahead approximation $\theta^{LA}$. Using the framework proposed by Powell [26], we use the following formulation:
@@
Optimized battery schedules are evaluated using a set of KPIs. The optimization targets the minimization of the equally weighted sum of the normalized electricity cost $C$ and carbon emissions $G$. The optimization score is defined as:
When grid-related Key Performance Indicators (KPIs) are included, the average includes the grid score $D$. The grid score is computed as the mean of the normalized ramping KPI $R$ and the Load Factor KPI $(1-L)$. The grid score, $D$, introduces switching costs in the optimization by accounting for the ramping costs associated with frequent changes in forecasting. Therefore, in order to provide a comprehensive analysis of the impact of switching costs on the optimization, in this study we consider both optimizations with and without the grid score. The average score with the grid cost is defined as:
[…]
@@ 2.2 Optimization
The optimal policy is computed for the current state of the system and the forecast for the next $H$ time steps. An online optimization algorithm typically determines the optimal actions $x_{t+1},...,x_{t+H}$ given a prediction window of length $H$. Every $v$ time steps, the optimization is rerun to generate a new plan. This algorithm is referred to as a Fixed Horizon Control (FHC) algorithm.
@@ 2.3 Forecast
Following the solution implemented in the winning submission of the competition, the forecasts are generated with gradient boosting regression trees and the linear least squares regression model. The gradient boosting implementation is LightGBM [27]. Predictions are made for the next 24 hours with 24 independent models. The scenario sets are generated with added gaussian noise. The details of the forecast models and the generation of the scenario sets are outlined in the competition paper by Nweye et al. [25] and the code repository of the winner solution is available on GitLab [28]. The evaluation conducted in this study is available in the code repository [29].
@@ 2.3.1 Rolling Origin Window
Figure 3: Vertical stability between rolling origin windows
in time-series forecast
The rolling horizon setup enables an update of the prediction using the most recent information. When the commitment period is shorter than the forecast horizon, the regular updates create overlaps between prediction windows, when the forecasts associated with a particular target time in the future are released in multiple revisions. The overlaps between windows with a rolling origin are shown in the Figure 3. The overlapping interval is highlighted in light grey.
@@ 2.3.2 Stability
As a potential factor influencing the performance of the policy, we are interested in the stability of the forecast. The stability of the forecast is defined as the consistency of the predictions over time. Following the categorization, first proposed in [18], we distinguish vertical and horizontal stability. The difference between the two concepts is illustrated in Figure 4.
@@ 3 Experimentation and Results
Further, we investigate the relationship between the performance of the FHC algorithm and properties of the forecast, namely accuracy and stability. In Figure 5, the accuracy and stability of the deterministic and stochastic forecasts are evaluated for different commitment periods. The comparison is made between the optimization with and without switching costs for the forecasts generated for out-of-sample buildings. Similar evaluation is done for the in-sample buildings and shown in A.4 It is observed from the plots that the forecast error generally decreases with shorter commitment periods, while the vertical stability is better with longer commitment periods. The horizontal stability demonstrates a trade-off where the optimal level is reached at a certain commitment period between 1 and 12 hours.
For stochastic forecast error, measured with the Energy Score, we observe an anomaly when the lowest error is observed for scenarios updated every 8 hours. The hypothesis is that this is due to the nature of the method of noise simulation, as outlined in [25]. The magnitude of added gaussian noise is proportional to the value of the respective point prediction, which produces low variation in the regions around zero, where this specific configuration captures most accurately, therefore the resulting scenario sets capture the real distribution with a low variation.
The plots of the optimization scores show different behavior depending on the presence of the switching costs. The optimization score without switching costs is optimal with shorter commitment periods, while the addition of switching costs adds more fluctuation to this relationship. The optimal commitment period remains at 1 for the point forecast, while the stochastic forecast enables an optimal performance at a longer commitment period.
@@ 4 Discussion
The addition of switching costs makes the algorithmic problem harder as it forces current actions to depend on beliefs about future cost functions. In Figure 6, we can see that the shape of the final load curve is significantly different when switching costs are included into the optimization. Without switching costs, the battery is charged and discharged while following the price and emissions cost signals. The resulting net load curve (in green) is compared to the perfect scenario (in yellow), where the battery is charged and discharged following the perfect forecast. In the case with switching costs, the optimal policy is regulating the net load curve to a smoother shape with less drops and peaks. The charging and discharging actions are the means of such regulation. However, due to the error in the forecast, the actual net load curve (in green) is not as smooth and flat as the optimal policy (in yellow). It generally follows the optimal policy, but with oscillation around it.
@@
Minimal commitment periods $v$ of the control algorithm, combined with imperfect predictions, can lead to a scenario where the control strategy is constantly “chasing” the optimal state based on imperfect prediction. This can aggravate the impact of switching costs, as the control actions may change frequently in response to new information, even when such changes are not cost-effective in the long run. While traditional forward-looking RHC formulations can include terms in the objective function to penalize switching costs, the longer commitment period provides a native mechanism that respects past actions of the FHC algorithm. This mechanism can reduce the need for frequent adjustments, thus avoiding some of the switching costs.
We identify two reasons for underperformance of the algorithm with a minimal commitment period (RHC, equivalent to FHC with $v_{O}=1$). Firstly, when the policy is computed for a period in a problem with switching cost, the past decisions are calculated as optimal with respect to future decisions given the forecast, however, after the policy is reviewed at the next iteration with an updated forecast, the optimality of the executed decision at the recent time step does not hold, since the future policies within the same horizon are reviewed. A simple illustrative example would be the battery is discharged first with an outlook for charging in the future when the price is more favorable. Once the forecast and policy updates, the updated policy is discharging again, however the battery has been discharged earlier. The second reason is that while the objective cost function does include the term for the switching costs $\beta||{x_{t}}-x_{t-1}||$, the switching costs are computed pairwise between loads at close time instances. However, since the total charging capacity of batteries is finite and limited, it is advantageous to regulate switching costs within longer action windows.
When prediction is inaccurate in the look-ahead window, the RHC is more effective than FHC with a commitment period. If we make predictions accurate for a small window $\gamma$ and then inaccurate for the remaining $(\omega-\gamma)$ steps of the lookahead window, the fixed horizon policy is affected by the inaccurate predictions whereas RHC only acts on the correct ones. This is typically the case for a multi-step horizon forecast when the nearest steps are predicted at a higher accuracy. The situation changes for a stochastic FHC, generally hedged better against prediction errors. We can see from the results that FHC with longer commitment period becomes viable when the optimization is computed for a set of scenarios, as opposed to a deterministic point forecast.
@@ 5 Conclusion
In conclusion, this study highlights the critical role that switching costs, forecast accuracy, and stability play in the design and operation of energy management systems. In summary, our analysis shows that in systems where forecasts guide decision making, stability of predictions can reduce the frequency and severity of policy revisions, thereby mitigating the financial and operational impacts of switching costs. Our findings have significant implications for the design of energy management systems. In the power system context, the switching costs are driven by imbalances, frequency deviations and additional stress on the power electronics. The current work suggests that enhancing the stability of forecasts can improve system performance by reducing the need for frequent adjustments and managing the trade-offs between immediate and future resource allocation more effectively. Furthermore, we propose $SDC$, a novel metric to evaluating the stability of probabilistic forecasts, which provides a robust measure of the dissimilarity between scenario sets.
Looking ahead, there are questions remaining to be addressed. For future research, we recommend further exploration into methods that enhance forecast stability without compromising accuracy. It would be beneficial to extend this work to other applications with switching costs, such as supply chain management and financial trading. Furthermore, from literature and current findings, there are indications that methods like AFHC and stochastic FHC, show more robustness in hedging against the forecast errors by aggregating policies with different underlying scenarios. However, they require more computational resources and are more complex to implement. An interesting avenue for future research would be defining the criteria for problems where these methods are most effective. These efforts would contribute to a deeper understanding of forecast-dependent optimization and improve the operational strategies of energy management systems facing an evolving energy landscape.

### v4

@@ 1 Introduction
To effectively navigate these challenges, it is beneficial to integrate forecasting of the problem environment with the optimization of decision-making processes. This integration, often referred to as the predict, then optimize approach, is prevalent in energy applications like electric vehicle charging [1, 2], battery scheduling [3, 4], and energy and flexibility dispatch problems [5, 6]. In many of these applications, decision-makers face switching costs — expenses incurred from updating operational plans. For example, transitioning operations from one state to another often involves costs due to physical limitations on ramping energy production up or down [7], managing server load balancing [8], or covering network fees associated with energy trading [9, 10].
Despite extensive research on online optimization with predictions, there remains a notable gap concerning how switching costs affect the synergy between forecasting models and optimization algorithms. This study addresses this gap by exploring the concepts of time-coupling and forecast stability within such systems. Time-coupling, where one decision directly affects subsequent decisions and system states, gives rise to switching costs. Forecast inaccuracies often lead to policy revisions, requiring adjustments to better align with the current system state and updated information. Striking a balance between forecast accuracy and the impact of switching costs is crucial. This paper explores how this balance shapes system design and informs operational strategies. Consequently, this research aims to answer the following questions:
@@ 1.1 Related work
The third category, Semi-direct methods, considers characteristics of the optimization problem but abstains from direct interaction during training of the forecast. These methods can take various forms, as potential improvements can occur at any stage outside of the training process. An example is found in the work by Kazmi and Paskevich [14], where those authors propose a Bayesian Optimization approach for integrating the downstream optimization problem into the hyperparameter tuning process of the forecast model. Other example is [3], where the authors propose iterating attention over training data statically and dynamically to improve the forecast model performance in the downstream task. Among the semi-direct methods, there is a noticeable gap in the literature regarding the operational aspects of integrating forecasting models and optimization deployment. A relevant study by Prat et al. [15] addresses the issue of determining the minimum forecast horizon for storage scheduling problems in a rolling-horizon approach. The study introduces a verifiable condition to check if a chosen planning horizon is sufficiently long. Other researchers study the effect of the frequency of forecast revision, which can be interpreted as both the frequency of the forecast model deployment and the frequency of model training. The frequency of retraining is studied by Spiliotis and Petropoulos [16], where the authors investigate different scenarios of updating the model fit for univariate exponential smoothing (ES) and univariate gradient boosting models (LightGBM).
We are particularly interested in studying the model deployment frequency. More specifically, we look at the applications of optimization with predictions in environments where switching costs are present. This topic has featured in research on online convex optimization (OCO) problems. In control theory, a widespread strategy for tackling online multi-step optimization challenges is the employment of the Receding Horizon Control (RHC) algorithm. Model Predictive Control (MPC) is the most common implementation of RHC that employs a dynamic model of the process and is widely used in online optimization. When rerun at every time step, this method commits to the first step of the future horizon while treating later decisions as advisory. The Fixed Horizon Control (FHC) algorithm is a generalization of the algorithm that commits to a fixed number of steps $v$ in the future before re-optimizing the policy. The difference between the two algorithms is illustrated in Figure 1. Throughout this paper, we will use the term FHC to align with the focus on commitment-specific optimization.
@@
Forecast stability has been recognized as a critical factor in planning. Building on the comprehensive framework by Godahewa et al. [20], which categorizes stability into vertical (variations across forecast updates for the same period) and horizontal (variations within a prediction window), this work investigates these dimensions in the context of energy management systems.
Unlike supply chain management, where freezing intervals [21, 22] and smoothing techniques [23] have been widely used to address system nervousness and the bullwhip effect [24], energy systems require more dynamic and frequent forecast updates. Switching costs are also observed in energy management systems, such as ramping costs for generating units [7, 25] or toggling server load balancing [8, 26]. However, the literature lacks a systematic exploration of how commitment levels influence the interplay between forecast stability and switching costs. By focusing on these aspects, the current research sets the stage for advancing the design and operation of energy management systems, particularly in applications involving energy storage optimization and real-time decision-making.
@@ 2.1 Problem statement
This paper addresses the challenge of optimizing battery storage management within a multi-agent energy management system. The setting of the problem is first outlined in the CityLearn Challenge 2022 [27] and based on a real-world dataset from a grid-interactive community. The CityLearn Challenge is an annual competition that focuses on the development of reinforcement learning agents for energy management in buildings. Participants are provided with a simulated environment that mimics real-world energy consumption and generation patterns. The goal is to optimize the operation of energy systems, such as batteries and HVAC systems, to minimize costs and emissions while maintaining comfort levels.
The series of CityLearn competitions, starting from 2020, was originally set up only for Reinforcement Learning (RL) solutions, where an agent learns to make decisions by performing actions and receiving feedback from the environment. However, from 2022 onwards, alternative methods such as Model Predictive Control (MPC), which uses a model to predict the future behavior of a system and optimize the control actions accordingly, and Rule-Based Control (RBC), which involves using predefined rules to make decisions based on the current state of the system, have also been employed. The competition provides a platform and the open-source simulation environment [^1] https://github.com/intelligent-environments-lab/CityLearn for implementing the operation of energy management systems with setups that include various energy assets, such as solar panels, batteries, heat pumps, connections to the grid, and electric vehicles.
The setup of the investigated problem is illustrated in Figure Figure 2. In general terms, we consider an online optimization problem with switching costs. The data of the competition has the hourly resolution. The core problem is multi-agent scheduling of energy storage over a period of a year. While the problem can be formulated similarly to RL, or a dynamic programming approach, in this paper, we particularly investigate the mechanics with a rolling-horizon approach. This means that the problem is solved iteratively over a planning horizon $H$. However, it is typical for only first $v$ steps to be implemented with the rest of the prediction horizon $H-v$ to be advisory, since the optimization is revised with a period of $v$ steps. The full details of the objective score and relevant constraints are provided in A.1. The problem includes finding an optimal state-dependent policy $X_{t}^{FHC}(S_{t}|\theta^{LA})$. The decision-maker is faced with a sequence of decisions, where the decision at each step is based on the current state $S_{t}$ of the system and the direct lookahead approximation $\theta^{LA}$. Using the framework proposed by Powell [28], we use the following formulation:
@@
Optimized battery schedules are evaluated using a set of Key Performance Indicators (KPIs). The optimization targets the minimization of the equally weighted sum of the normalized electricity cost $C$ and carbon emissions $G$. When grid-related metrics are included, the average includes the grid score $D$, which is computed as the mean of the normalized ramping KPI $R$ and the Load Factor KPI $(1-L)$. The grid score, $D$, introduces switching costs in the optimization by accounting for the ramping costs associated with frequent changes in forecasting. Therefore, in order to provide a comprehensive analysis of the impact of switching costs on the optimization, the combined optimization scores are defined with and without the grid score:
Average Score $=\operatorname{avg}\left(\frac{C_{\text{entry}}}{C_{\text{no\ battery}}},\frac{G_{\text{entry}}}{G_{\text{no\ battery}}}\right)$ (2)
Average Score (with switching costs) $=\operatorname{avg}\left(\frac{C_{\text{entry}}}{C_{\text{no\ battery}}},\frac{G_{\text{entry}}}{G_{\text{no\ battery}}},D\right)$ (3)
@@ 2.2 Optimization
The problem is solved using the Fixed Horizon Control (FHC) algorithm, which is a configuration of model predictive control (MPC) with a fixed commitment horizon. The optimal policy is computed for the current state of the system and the forecast for the next $H$ time steps. An online optimization algorithm typically determines the optimal actions $x_{t+1},...,x_{t+H}$ given a prediction window of length $H$. Every $v$ time steps, the optimization is rerun to generate a new plan. This algorithm is referred to as a Fixed Horizon Control (FHC) algorithm.
@@ 2.3 Forecast
The target variable for forecasting is the aggregate uncontrollable load of the energy community in Figure 2. Following the solution implemented in the winning submission of the competition, the forecasts are generated with gradient boosting regression trees and the linear least squares regression model. The gradient boosting implementation is LightGBM [29]. Predictions are made directly for the next 24 hours with 24 independent models. The scenario sets are generated with added gaussian noise. The details of the forecast models and the generation of the scenario sets are outlined in the competition paper by Nweye et al. [27] and the code repository of the winner solution is available on GitLab [30]. The evaluation conducted in this study is available in the code repository [31].
@@ 2.3.1 Rolling Origin Window
Figure 3: Updating Rolling Origin Window. The forecast
is updated every $v_{f}$ time steps and possible overlaps between prediction windows are shown.
The rolling horizon setup enables an update of the prediction using the most recent information. When the commitment period is shorter than the forecast horizon, the regular updates create overlaps between prediction windows, when the forecasts associated with a particular target time in the future are released in multiple revisions. The resulting overlapping interval is highlighted in light grey in Figure 3. These overlaps are considered when evaluating stability of the forecast.
@@ 2.3.2 Stability
Stability of the forecast is defined as the consistency of the predictions over time. Following the categorization, first proposed in [20], we distinguish vertical and horizontal stability. The difference between the two concepts is illustrated in Figure 4.
@@ 3 Experimentation and Results
Further, we investigate the relationship between the performance of the FHC algorithm and properties of the forecast, namely accuracy and stability. In Figure 5, the accuracy and stability of the deterministic and stochastic forecasts are evaluated for different commitment periods between 1 and 12. The comparison is made between the optimization with and without switching costs for the forecasts generated for out-of-sample buildings. Similar evaluation is done for the in-sample buildings and shown in A.4. It is observed from the plots that the forecast error generally decreases with shorter commitment periods, while the vertical stability is better with longer commitment periods. The horizontal stability demonstrates a trade-off where the optimal level is reached at a certain commitment period between 1 and 12 hours. For stochastic forecast error, measured with the Energy Score, we observe an anomaly when the lowest error is observed for scenarios updated every 8 hours. The hypothesis is that this is due to the nature of the method of noise simulation, as outlined in [27]. The magnitude of added gaussian noise is proportional to the value of the respective point prediction, which produces low variation in the regions around zero, where this specific configuration captures most accurately, therefore the resulting scenario sets capture the real distribution with a low variation.
The plots of the optimization scores show different behavior depending on the presence of the switching costs. The optimization score without switching costs is optimal with shorter commitment periods, while the addition of switching costs adds more fluctuation to this relationship. The optimal commitment period remains at 1 for the point forecast, while the stochastic forecast enables an optimal performance at a longer commitment period. This observation is notable as the most expected behavior is that best performance is achieved with shortest revision periods for both forecast types, as more up-to-date forecasts are provided.
@@ 4 Discussion
The addition of switching costs makes the algorithmic problem harder as it forces current actions to depend on beliefs about future cost functions. In Figure 6, we can see that the shape of the optimal EMS net load curve is significantly different when switching costs are included into the optimization. Without switching costs, the battery is charged and discharged while following the price and emissions cost signals, indicated in the bottom subplot. The plot also highlights the difference in net load between the storage management with oracle information (in yellow) and forecast-based management (in blue). The oracle information, also referred to as the perfect forecast, is the ideal scenario where the EMS uses the perfect knowledge of the future load signals. Without considering switching costs, it is observed that scheduling with the oracle information has higher spikes. We assume that the oracle information given optimizing agent the confidence to charge at maximum capacity when the price and emissions are low. However, the forecast-based management is more conservative in its actions, as it is uncertain about the future load signals.
The figure illustrates the difference in net load for two scenarios: one without switching costs and one with. When switching costs are considered, the optimal policy smooths the net load curve, resulting in a more stable profile with fewer fluctuations. The charging and discharging actions contribute to this smoothing, making the aggregate net load more stable. However, the presence of forecast errors exacerbates the switching costs. The aggregate EMS-controlled net load curve (in green) is not as smooth and flat as the optimal policy (in yellow). While it generally follows the optimal policy, it oscillates around it.
@@
Minimal commitment periods $v$ of the control algorithm, combined with imperfect predictions, can lead to a scenario where the control strategy is constantly “chasing” the optimal state based on imperfect prediction. This can aggravate the impact of switching costs, as the control actions may change frequently in response to new information, even when such changes are not cost-effective in the long run. In line with the theoretical analysis in Appendix A.3, the upper bound on the competitive difference of the FHC algorithm is dependent on the trade-off between switching costs and forecast errors.
While the MPC objective functions factors in the switching costs, we identify two factors that can lead to underperformance of FHC with low commitment period ($v_{O}=1$), particularly when probabilistic forecasts are used. Firstly, when the policy is computed for a finite period in a problem with switching cost, the past decisions are calculated as optimal with respect to future decisions given the forecast, however, after the policy is reviewed at the next iteration with an updated forecast, the optimality of the previously executed decision may be suboptimal given the updated information. A simple illustrative example would be the battery is discharged first with an outlook for charging in the future when the price is more favorable. Once the forecast and policy updates, the updated policy is discharging again, however the battery has been discharged earlier. The second reason is that while the objective cost function does include the term for the switching costs $\beta||{x_{t}}-x_{t-1}||$, the switching costs are computed pairwise between loads at close time instances. However, since the total charging capacity of batteries is finite and limited, it is advantageous to regulate switching costs within longer action windows.
Other important observation is that the FHC algorithm with longer commitment periods only becomes viable when the optimization is computed for a set of scenarios, as we can see from the results in Figure 5. It is hypothesized, that the aggregation of policies over different scenarios provides a more robust policy that hedges against the forecast errors. Generally, within a forecast horizon, the nearest steps are predicted with higher accuracy than further steps. Longer commitment periods aggravate the impact of forecast errors due to time coupling, as the suboptimal decisions made in the past are amplified in the future. A stochastic FHC algorithm is better equipped to handle errors in the forecast, as it aggregates policies over different scenarios, therefore, the impact of forecast errors is mitigated and longer commitment periods are more viable.
@@ 5 Conclusion
To summarize, the inclusion of switching costs in the FHC algorithm introduces a trade-off between forecast errors and switching costs. While traditional FHC formulations can penalize switching costs in the objective function, longer commitment periods provide a natural mechanism to use a stable forecast and reduce the need for frequent adjustments. This approach is more viable when optimization is computed over a set of scenarios, since the aggregation of policies offers a more robust policy, hedging against increasing forecast errors in a forecast horizon.
In conclusion, this study highlights the critical role that switching costs, forecast accuracy, and stability play in the design and operation of energy management systems. Our analysis shows that in systems where forecasts guide decision making, stability of predictions can reduce the frequency and switching costs of policy revisions, thereby mitigating the financial and operational impacts. Our findings have significant implications for the design of energy management systems. In the power system context, the switching costs are driven by imbalances, frequency deviations and additional stress on the power electronics. The current work suggests that enhancing the stability of forecasts leads to stability of policy which can improve system performance by managing the trade-offs between forecast accuracy and switching costs. Furthermore, we propose $SDC$, a novel metric to evaluating the stability of probabilistic forecasts, which provides a robust measure of the dissimilarity between scenario sets.
Looking ahead, there are questions remaining to be addressed. For future research, we recommend further exploration into methods that enhance forecast stability without compromising accuracy. Further research into the link between forecast quality and optimization performance is needed. Parameters such as forecast horizon, commitment period, and the number of scenarios are subject to a study for deeper understanding of their impact on the optimization performance and dependencies to the forecast accuracy and stability. It would be beneficial to extend this work to other applications with switching costs, such as supply chain management and financial trading. From literature and current findings, methods like AFHC and stochastic FHC, show more robustness in hedging against the forecast errors, however, they require more computational resources and are more complex to implement. An interesting avenue for future research would be defining the criteria for problems where these methods are most effective. These efforts would contribute to a deeper understanding of forecast-dependent optimization and improve the operational strategies of energy management systems facing an evolving energy landscape.